OpenStudy (cenaida):

Solve the following system by elimination. 3x - 2y = 18 5x + 2y = 14 A. (0, 7) B. (4, -3) C. (6, -6) D. (2, 2)

OpenStudy (cenaida):

@abb0t is it B?

OpenStudy (aravindg):

How did you get B?

OpenStudy (abb0t):

Plug in those values for \(b\). Does it equal 18 or 14 when you plug them in?

OpenStudy (kenljw):

3x - 2y = 18 5x + 2y = 14 add 8x = 32 x = 4 place into either equation to get y 12 - 2y = 18 -2y = 6 y = -3 QED